OpenStudy (anonymous):

Solve by using the perfect squares method x^2-12x+36=0

OpenStudy (lieutenantgeneral):

x = -6 (x+6)(x +6) = 0 x2+12x+36=0 x2+6x+6x+36=0 (x+6)(x+6)=0 x=-6